# Concept – An Evaluation Protocol on Conversation Recommender Systems with System- and User-centric Factors

### v1

@@ Appendix C Characteristics Categorization
We consolidate both system- and user-centric factors as the starting point. Interdisciplinary research on conversational AI Chaves and Gerosa (2021); Reeves and Nass (1996); Fogg (2003) groups its characteristics that impact the user experience into three types. Concept then tailors them for CRS, which conceptualizes expected characteristics into six primary abilities for the inclusive CRS evaluation. In Table 8, we elaborate the inclusive CRS evaluation protocol, with the special focus on the definitions of each abilities and skills, and how we evaluate them.
@@
C.1 Recommendation Intelligence
It requires CRS to learn from conversations and evolve toward making recommendations as the conversation advances. Taxonomy in Chaves and Gerosa (2021) highlights conversational intelligence, which pertains to the AI’s ability to facilitate conversations [^11] We evaluate CRS’s conversational proficiency within the context of social intelligence.. We highlight CRS’s capacity to model user preferences from conversations and create recommendations. As a basic function from system-centric perspectives, CRS needs to produce high-quality and reliable recommendations to fit user preferences based on the conversation history Chen et al. (2019); Ma et al. (2020); Zhou et al. (2021). In this regard, our categorization encompasses two primary abilities.
Quality. As the basic requirements, CRS should provide precise recommendations using minimal conversation turns, which are the crucial aspects that influence user satisfaction Siro et al. (2023); Gao et al. (2021). Apart from evaluating the quality of items recommended by the recommendation module and the conversation module [^12] Conversation module can conduct recommendation Chen et al. (2019), motivated us to gauge its quality. separately, we emphasize the importance of user acceptance rate, reflecting the practical effectiveness and relevance of the recommendations in practice.
Specially, we consider computational metrics for automatic evaluation following previous works Wang et al. (2023d); Zhang et al. (2023); Yu et al. (2023), including $Recall@k(k=1,10,25,50)$, recommendation success rate $(SR@k,k=3,5,10)$, user acceptance rate ($AR$), and average turns ($AT$) required to reach the successful recommendation. Note that we determine user acceptance of the recommendation by checking for the presence of the special token ’[END]’ in the user’s response.
Reliability. CRS should deliver robust and consistent recommendations that account for contextual nuances. In practical situations, users with similar preferences may express themselves differently. It would be frustrating to recommend divergent items for two semantically similar inputs Oh et al. (2022). This not only diminishes the user experience but can also be disruptive or even perilous for critical applications in healthcare and finance Tran et al. (2021); Zibriczky12 (2016). In less critical scenarios, if two recommended items are inconsistent but align with user preferences, they are often viewed as diverse recommendations, creating more opportunities for relevant but less popular items Yang et al. (2021). We also consider this when evaluating.
@@ C.2 Social Intelligence
It requires CRS to produce adequate social behavior for the recommendation during the conversations. Taxonomy in Chaves and Gerosa (2021) highlights social intelligence, which focuses on habitual social protocols. As evidenced by the Media Equation Theory Reeves and Nass (1996); Fogg (2003), users tend to engage with the machine in a manner that mirrors person-to-person conversations. They have high expectations for the machine to adhere to the cooperative principle Nass et al. (1996), and instinctively expect the machine to demonstrate empathy Björkqvist et al. (2000); Hayati et al. (2020) in order to facilitate comfortable conversations in social situations. These highlight the need for CRS to act cooperatively and be aware of the user’s social needs during the conversation, facilitating the design of CRS with perceived humanness Neururer et al. (2018); Jacquet et al. (2018, 2019). In this regard, our categorization encompasses two primary abilities, including cooperation and social awareness.
@@
Cooperation. CRS should follow the cooperative principle to achieve effective and comfortable conversations in common social situations Reeves and Nass (1996); Fogg (2003); Chaves and Gerosa (2021). The cooperative principle guides individuals to communicate in a way that is helpful and contributes to the overall success of the conversation Nass et al. (1996). This is achieved by following four “Maxims [^13] Maxim is a phrase or saying that includes a rule or moral principle about how one should behave. of Conversation" Grice (1975, 1989); SARI (2020), which serve as a foundation for refining the cooperative ability of CRS: 1) Manner. CRS should respond in a manner that is easily understood and clearly expressed. 2) Sincerity. CRS should communicate sincerely, without deception or pretense, and ensure that its responses are backed by sufficient evidence. 3) Response Quality. CRS should provide the necessary level of information for the conversation without overwhelming the user with unnecessary details. 4) Relevance. CRS should ensure that its responses contribute to identifying user preferences and making recommendations.
@@
Social Awareness. CRS needs to meet users’ social expectations in practice, showing care, empathy, and establishing rapport with them Björkqvist et al. (2000). This facilitates the authenticity of CRS Neururer et al. (2018). To achieve this, a recent study Hayati et al. (2020) identified eight social strategies for building rapport with users. For instance, CRS could engage in self-disclosure and share its subjective opinion about a movie to establish a social connection with users.
@@
As for the evaluation implementations, We mainly resort to an LLM-based evaluator with elaborately designed instructions and ability-specific score rubrics (ranging from $1-5$) for evaluation (cf. Appendix A.2.3). As for the sincerity evaluation, we value the ratio of non-existent items and deceptive tactics. For non-existent items, we have tallied how many items recommended by the CRS do not exist in the dataset. As for the deceptive tactics, we focus on the items accepted by users and tally how many of these do not align with the users’ pre-defined preferences. If users accept these misleading items, we consider the CRS to be employing a deceptive tactics, leading users to believe that the recommended items meet their preferences.
@@
C.3 Personification
It requires CRS to perceive the identity of itself and the personality representation of users. Taxonomy in Chaves and Gerosa (2021) highlights personification of conversational AI in terms of identity and personality, which value the different traits of conversational AI. This requires CRS to have self-awareness of its role as a conversational recommendation system, distinguishing itself from sales systems or other types of systems. This helps CRS operate within its intended scope and purpose, ensuring that it effectively fulfills its role by providing persuasive and honest explanations to encourage acceptance Jannach et al. (2021); Zhou et al. (2022) for diverse users Thompson et al. (2004). In this regard, our categorization encompasses two primary abilities.
Identity. CRS should provide persuasive and honest explanations to boost the acceptance of the recommendations. In real-world situations, users often have little knowledge about the recommended items. A persuasive explanation has the potential to alter a user’s perception of the recommendation, consequently increasing its acceptance Jannach et al. (2021); Zhou et al. (2022); McSherry (2005). However, generating explanations must be handled with great care to ensure adequacy and avoid resorting to misleading or deceptive strategies (i.e., sales pitches with deceptive tactics), as this could diminish user trust and loyalty towards the CRS Gkika and Lekakos (2014). More importantly, deceptive strategies also violate the maxim of sincerity required by the cooperation ability (cf. Section C.2). We emphasize the importance of integrating ethical guidelines into CRS and promoting an ethical CRS that mitigates any potential deceptive behavior.
@@ C.3 Personification
Coordination. CRS should be proficient in serving users possessing various personas without prior coordination. Attributing personality to a conversational AI ensures that its behaviors stand in agreement with the users’ expectations in a particular context Chaves and Gerosa (2021). In the case of CRS, it is pervasive for CRS to encounter users with varying personas, such as preferences and social communication patterns following the deployment of CRS in real-world scenarios Thompson et al. (2004). This requires CRS to demonstrate different personalities and adjust its behavior to align with situational, emotional context, and users’ preferences Katayama et al. (2019); Svikhnushina et al. (2021); Svikhnushina (2023); Zhou et al. (2020). A key measurement is to proficiently engage and serve users without prior coordination Mirsky et al. (2022).
@@
Figure 7: Evaluation of social-centric characteristics across each benchmark dataset
Appendix D Reliability of Concept
We have established fixed values for the temperature and seed parameters to guarantee replicability. We show that Concept scoring is unbiased and aligns with the results of human evaluations.
D.1 Bias Analysis
LLM-based evaluation is prone to biases Wang et al. (2023e), but can be eliminated by using fine-grained scoring rubrics Ye et al. (2023); Wang et al. (2023d); Liu et al. (2023d). We found that this conclusion also holds true in our case.
  • Length bias. As in Figure 8, Concept is robust to the length bias [^14] LLMs prefer longer responses Wu and Aji (2023).. Although CHATCRS tends to generate long responses, this does not mean that longer responses will result in higher scores.
  • Self-enhancement bias. Regarding the self-enhancement bias [^15] Bias of CHATGPT to favor high scores for its generated content Wang et al. (2023c)., our human evaluation revealed that this bias does not exist. The content produced by CHATCRS is notably superior to that of other CRS models, which struggle to maintain coherent conversations and at times produce incomprehensible content. For instance, KBRD generated the following: "This movie is about a man who’s a woman who gets murdered and the world." To quantitatively measure this, in the generated conversation data, the replies from the user simulator for all CRS models are sourced from ChatGPT. Therefore, we do not evaluate the content of the user simulator’s responses, but solely focus on the responses from different CRS models. In Appendix D.2, our human evaluation revealed that the evaluation results of Concept closely align with human evaluations. Additionally, we randomly selected conversation data from three other models and contrasted them with the conversation data of CHATCRS utilized in the human evaluation experiment, and observed that the response content of CHATCRS consistently surpasses that of the other models.
Figure 8: Length bias evaluation. We require CHATGPT to provide an overall performance score based on all the ability-specific scores (i.e., Y-axis) and then plot the relationship between various CRS reply lengths (X-axis) and the scores. Although CHATCRS tends to generate long responses, this does not mean that longer responses will result in higher scores.
Figure 9: Evaluation results on Human Likeness, overall performance, and user satisfaction.
@@ D.1 Bias Analysis
D.2 Human Evaluation
Previous findings highlight that offering detailed scoring rubrics or criteria contributes to achieving consistent and aligned evaluations with human assessments Liu et al. (2023d); Wang et al. (2023d), ultimately resulting in a dependable LLM-based evaluator as a viable alternative to human evaluators. We also conduct human evaluation to analyze the consistency. Our human evaluation is used to examine the capability of the CHATGPT-based user simulator being a movie seeker, while checking the correlation between the CHATGPT-based evaluation results and human evaluation results. The results are as follows, which indicate the reliability of our Concept.
Human evaluation results based on the conversation data and the LLM-based evaluation results are highly related, with a correlation coefficient of 61.24% and Krippendorff’s alpha of 53.10%. Due to budget limitations, we are unable to hire a large number of individuals with various personas and age groups for human evaluation. Instead, our human evaluation is conducted by two human evaluators. In particular, we randomly selected 120 conversations involving CHATCRS and users with various personas and age groups. Given the numerous scoring aspects in our Concept, it is prone to inconsistent ratings across different aspects. To address this, we had ChatGPT score all aspects of these 120 conversations and provide an overall performance score. Subsequently, each human only needed to provide an overall score, without having to rate all aspects. During the scoring process, each evaluator independently scored these 120 conversations based on our detailed criteria. They then discussed any disagreements. We calculated the inter-annotator reliability using Krippendorff’s alpha, achieving 41.34%. To ensure consistency and enhance the robustness of the evaluations, we took the average of the scores from both evaluators as the final result for each conversation. We then compared the human evaluation results with those from an LLM-based evaluator and found a correlation of 61.24% and Krippendorff’s alpha of 53.10%. This indicates the reliability of our LLM-based evaluator, which is consistent with previous findings Wang et al. (2023d).
@@ D.2 Human Evaluation
Appendix E Additional Analysis
@@
E.1 Overall Performance
We report more results in Figure 9 in terms of the Human Likeness, overall performance, and user satisfaction. This is achieved by prompting the LLM-based evaluator using the detailed results of all ability-specific scores and fine-grained scoring rubrics. Refer Appendix F for details on prompts.
E.2 Fine-grained Analysis on each Benchmark Dataset

### v2

@@ Appendix D Characteristics Categorization
We consolidate both system- and user-centric factors as the starting point. Interdisciplinary research on conversational AI Chaves and Gerosa [2021], Reeves and Nass [1996], Fogg [2003] groups its characteristics that impact the user experience into three types. Concept then tailors them for CRS, which conceptualizes expected characteristics into six primary abilities for the inclusive CRS evaluation. In Table 8, we elaborate the inclusive CRS evaluation protocol, with the special focus on the definitions of each abilities and skills, and how we evaluate them.
@@
D.1 Recommendation Intelligence
It requires CRS to learn from conversations and evolve toward making recommendations as the conversation advances. Taxonomy in Chaves and Gerosa [2021] highlights conversational intelligence, which pertains to the AI’s ability to facilitate conversations [^12] We evaluate CRS’s conversational proficiency within the context of social intelligence.. We highlight CRS’s capacity to model user preferences from conversations and create recommendations. As a basic function from system-centric perspectives, CRS needs to produce high-quality and reliable recommendations to fit user preferences based on the conversation history Chen et al. [2019], Ma et al. [2020], Zhou et al. [2021]. In this regard, our categorization encompasses two primary abilities.
Quality. As the basic requirements, CRS should provide precise recommendations using minimal conversation turns, which are the crucial aspects that influence user satisfaction Siro et al. [2023], Gao et al. [2021]. Apart from evaluating the quality of items recommended by the recommendation module and the conversation module [^13] Conversation module can conduct recommendation Chen et al. [2019], motivated us to gauge its quality. separately, we emphasize the importance of user acceptance rate, reflecting the practical effectiveness and relevance of the recommendations in practice.
Specially, we consider computational metrics for automatic evaluation following previous works Wang et al. [2023d], Zhang et al. [2023], Yu et al. [2023], including $Recall@k(k=1,10,25,50)$, recommendation success rate $(SR@k,k=3,5,10)$, user acceptance rate ($AR$), and average turns ($AT$) required to reach the successful recommendation. Note that we determine user acceptance of the recommendation by checking for the presence of the special token ’[END]’ in the user’s response.
Reliability. CRS should deliver robust and consistent recommendations that account for contextual nuances. In practical situations, users with similar preferences may express themselves differently. It would be frustrating to recommend divergent items for two semantically similar inputs Oh et al. [2022]. This not only diminishes the user experience but can also be disruptive or even perilous for critical applications in healthcare and finance Tran et al. [2021], Zibriczky12 [2016]. In less critical scenarios, if two recommended items are inconsistent but align with user preferences, they are often viewed as diverse recommendations, creating more opportunities for relevant but less popular items Yang et al. [2021]. We also consider this when evaluating.
@@ D.2 Social Intelligence
It requires CRS to produce adequate social behavior for the recommendation during the conversations. Taxonomy in Chaves and Gerosa [2021] highlights social intelligence, which focuses on habitual social protocols. As evidenced by the Media Equation Theory Reeves and Nass [1996], Fogg [2003], users tend to engage with the machine in a manner that mirrors person-to-person conversations. They have high expectations for the machine to adhere to the cooperative principle Nass et al. [1996], and instinctively expect the machine to demonstrate empathy Björkqvist et al. [2000], Hayati et al. [2020] in order to facilitate comfortable conversations in social situations. These highlight the need for CRS to act cooperatively and be aware of the user’s social needs during the conversation, facilitating the design of CRS with perceived humanness Neururer et al. [2018], Jacquet et al. [2018, 2019]. In this regard, our categorization encompasses two primary abilities, including cooperation and social awareness.
@@
Cooperation. CRS should follow the cooperative principle to achieve effective and comfortable conversations in common social situations Reeves and Nass [1996], Fogg [2003], Chaves and Gerosa [2021]. The cooperative principle guides individuals to communicate in a way that is helpful and contributes to the overall success of the conversation Nass et al. [1996]. This is achieved by following four “Maxims [^14] Maxim is a phrase or saying that includes a rule or moral principle about how one should behave. of Conversation" Grice [1975, 1989], SARI [2020], which serve as a foundation for refining the cooperative ability of CRS: 1) Manner. CRS should respond in a manner that is easily understood and clearly expressed. 2) Sincerity. CRS should communicate sincerely, without deception or pretense, and ensure that its responses are backed by sufficient evidence. 3) Response Quality. CRS should provide the necessary level of information for the conversation without overwhelming the user with unnecessary details. 4) Relevance. CRS should ensure that its responses contribute to identifying user preferences and making recommendations.
@@
Social Awareness. CRS needs to meet users’ social expectations in practice, showing care, empathy, and establishing rapport with them Björkqvist et al. [2000]. This facilitates the authenticity of CRS Neururer et al. [2018]. To achieve this, a recent study Hayati et al. [2020] identified eight social strategies for building rapport with users. For instance, CRS could engage in self-disclosure and share its subjective opinion about a movie to establish a social connection with users.
@@
As for the evaluation implementations, We mainly resort to an LLM-based evaluator with elaborately designed instructions and ability-specific score rubrics (ranging from $1-5$) for evaluation (cf. Appendix B.2.3). As for the sincerity evaluation, we value the ratio of non-existent items and deceptive tactics. For non-existent items, we have tallied how many items recommended by the CRS do not exist in the dataset. As for the deceptive tactics, we focus on the items accepted by users and tally how many of these do not align with the users’ pre-defined preferences. If users accept these misleading items, we consider the CRS to be employing a deceptive tactics, leading users to believe that the recommended items meet their preferences.
@@
D.3 Personification
It requires CRS to perceive the identity of itself and the personality representation of users. Taxonomy in Chaves and Gerosa [2021] highlights personification of conversational AI in terms of identity and personality, which value the different traits of conversational AI. This requires CRS to have self-awareness of its role as a conversational recommendation system, distinguishing itself from sales systems or other types of systems. This helps CRS operate within its intended scope and purpose, ensuring that it effectively fulfills its role by providing persuasive and honest explanations to encourage acceptance Jannach et al. [2021], Zhou et al. [2022] for diverse users Thompson et al. [2004]. In this regard, our categorization encompasses two primary abilities.
Identity. CRS should provide persuasive and honest explanations to boost the acceptance of the recommendations. In real-world situations, users often have little knowledge about the recommended items. A persuasive explanation has the potential to alter a user’s perception of the recommendation, consequently increasing its acceptance Jannach et al. [2021], Zhou et al. [2022], McSherry [2005]. However, generating explanations must be handled with great care to ensure adequacy and avoid resorting to misleading or deceptive strategies (i.e., sales pitches with deceptive tactics), as this could diminish user trust and loyalty towards the CRS Gkika and Lekakos [2014]. More importantly, deceptive strategies also violate the maxim of sincerity required by the cooperation ability (cf. Section D.2). We emphasize the importance of integrating ethical guidelines into CRS and promoting an ethical CRS that mitigates any potential deceptive behavior.
@@ D.3 Personification
Coordination. CRS should be proficient in serving users possessing various personas without prior coordination. Attributing personality to a conversational AI ensures that its behaviors stand in agreement with the users’ expectations in a particular context Chaves and Gerosa [2021]. In the case of CRS, it is pervasive for CRS to encounter users with varying personas, such as preferences and social communication patterns following the deployment of CRS in real-world scenarios Thompson et al. [2004]. This requires CRS to demonstrate different personalities and adjust its behavior to align with situational, emotional context, and users’ preferences Katayama et al. [2019], Svikhnushina et al. [2021], Svikhnushina [2023], Zhou et al. [2020]. A key measurement is to proficiently engage and serve users without prior coordination Mirsky et al. [2022].
@@
Appendix E Human Evaluation
Previous findings highlight that offering detailed scoring rubrics or criteria contributes to achieving consistent and aligned evaluations with human assessments Liu et al. [2023d], Wang et al. [2023d], ultimately resulting in a dependable LLM-based evaluator as a viable alternative to human evaluators. We also conduct human evaluation to analyze the consistency. Our human evaluation is used to examine the capability of the CHATGPT-based user simulator being a movie seeker, while checking the correlation between the CHATGPT-based evaluation results and human evaluation results. The results are as follows, which indicate the reliability of our Concept.
Human evaluation results based on the conversation data and the LLM-based evaluation results are highly related, with a correlation coefficient of 61.24% and Krippendorff’s alpha of 53.10%. Due to budget limitations, we are unable to hire a large number of individuals with various personas and age groups for human evaluation. Instead, our human evaluation is conducted by two human evaluators. In particular, we randomly selected 120 conversations involving CHATCRS and users with various personas and age groups. Given the numerous scoring aspects in our Concept, it is prone to inconsistent ratings across different aspects. To address this, we had ChatGPT score all aspects of these 120 conversations and provide an overall performance score. Subsequently, each human only needed to provide an overall score, without having to rate all aspects. During the scoring process, each evaluator independently scored these 120 conversations based on our detailed criteria. They then discussed any disagreements. We calculated the inter-annotator reliability using Krippendorff’s alpha, achieving 41.34%. To ensure consistency and enhance the robustness of the evaluations, we took the average of the scores from both evaluators as the final result for each conversation. We then compared the human evaluation results with those from an LLM-based evaluator and found a correlation of 61.24% and Krippendorff’s alpha of 53.10%. This indicates the reliability of our LLM-based evaluator, which is consistent with previous findings Wang et al. [2023d].
Figure 9: Human Likeness, overall performance, and user satisfaction.
@@ Appendix E Human Evaluation
Appendix F Additional Analysis
@@
F.1 Overall Performance
We report more results in Figure 9 in terms of the Human Likeness, overall performance, and user satisfaction. This is achieved by prompting the LLM-based evaluator using the detailed results of all ability-specific scores and fine-grained scoring rubrics. Refer Appendix G for details on prompts.
F.2 Fine-grained Analysis on each Benchmark Dataset
Figure 10: Evaluation of social-centric characteristics